# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ 4.1 Autocorrelation of the noise
Figure 1: Autocorrelations of the SGD noise observed over a span of 20 epochs, equivalent to 20,000 update steps. This data is collected from a later phase in the training process. The autocorrelation is projected onto 5,000 Hessian eigenvectors, and the result is averaged. The theoretical prediction Equation 2 is also displayed along with a $2\sigma$-interval, where $\sigma$ represents the expected standard deviation of the SGD noise. The zero-point correlation is omitted as it is inherently equal to one.
Autocorrelation of the noise means we are considering the correlation between two noise terms, but they stem from different update steps. In the case of SGD, when we sample the examples without replacement while keeping the weight vector $\bm{\mathbf{\theta}}$ constant, there are inherent anti-correlations in the noise. Here, we provide a brief conceptual motivation followed by the exact result where we direct the reader to Appendix E for a comprehensive derivation.
Suppose the ratio $M\coloneqq N/S$ is an integer, where $N$ represents the total number of examples, and $S$ is the batch size. This condition allows for the even distribution of examples among batches, ensuring each example is presented exactly once per epoch. Given a fixed weight vector $\bm{\mathbf{\theta}}$ and computing every gradient $\bm{\mathbf{g}}_{k}(\vec{\theta})$ for one epoch, it is evident that their mean equates to the total gradient $\bm{\mathbf{\nabla}}L(\bm{\mathbf{\theta}})$. Consequently, the sum of the noise components $\bm{\mathbf{\delta g}}_{k}(\bm{\mathbf{\theta}})=\vec{g}_{k}(\vec{\theta})-\vec{\nabla}L(\vec{\theta})$ for one epoch must be zero. This condition implies that two noise components within the same epoch are inherently anti-correlated.
This anti-correlation between two noise terms of the same epoch can therefore be expressed in terms of the the equal-time correlation, multiplied by a factor of $\frac{1}{M-1}$ with $M$ being the number of batches per epoch. Next, we consider the probability that two batches $k$ and $k+h$ with an update step difference $h$ are from the same training epoch. This probability is given by the factor $\frac{M-|h|}{M}$ for $|h|\leq M$, and zero otherwise. By combining with the above anti-correlation, we arrive at the correlation formula:
@@
If the total number of examples $N$ is an integer multiple of the batch size $S$ and the parameters $\bm{\mathbf{\theta}}$ of a network are kept fixed, then the autocorrelation formula for the gradient noise of an epoch-based learning schedule, where the examples for one epoch are drawn without replacement, is given by
${\rm cov}\bigl{(}\vec{\delta g}_{k}(\vec{\theta}),\vec{\delta g}_{k+h}(\vec{\theta})\bigr{)}=$
$\bm{\mathbf{C}}(\vec{\theta})\cdot\left(\delta_{h,0}-\mathbf{1}_{\{1,...,M\}}(|h|)\frac{M-|h|}{M(M-1)}\right)\ ,$ (2)
@@
In the above theorem $\mathbf{1}_{A}(k)$ represents the indicator function over the set A, which is one for $k\in A$ and zero otherwise and $\delta_{i,j}$ represents the Kronecker delta. The actual noise autocorrelation is illustrated in Figure 1, with the experimental details elaborated in Section 5.2. The complete calculation is available in Appendix E, highlighting the relationship $\bm{\mathbf{C}}(\vec{\theta})=(1/S)\cdot(1-S/N)\cdot\bm{\mathbf{C_{0}}}(\vec{\theta})$ with the gradient sample covariance matrix $\bm{\mathbf{C_{0}}}$.
It is important to note that the above formula is only applicable for a static weight vector. During training, weights alter with each update step, which could potentially modify this relationship. Nevertheless, if the weights remain relatively consistent over one epoch, this correlation approximately holds true. In a simple quadratic model one can also construct a noise which would be independent of the weight vector (see Appendix B), for which then this correlation holds true without restrictions on the weights. Either way, for the later stages of training, such a noise constancy assumption seems to be a good approximation, as evidenced by the close fit of the data to the theory in Figure 1.
Further numerical investigations show, that even in the case where $N$ is not an integer multiple of $S$, the later deduced relations for the variances still hold (see Appendix J).
4.2 Correlation time definition
@@
To gain a more comprehensive understanding of the weight variance behavior, we additionally scrutinize the velocity variance and the ratio between them. We label this ratio, scaled by a factor of two, as the correlation time $\tau_{i}$ (see Equation 3). This definition aligns with that of the velocity correlation time, hence justifying the nomenclature. The equivalence stands under general assumptions that are satisfied in our problem setup described in Section 4.3 (see Section C.4).
The assumptions encompass: (i) Existence and finiteness of $\bm{\mathbf{\Sigma}}\coloneqq$ cov$(\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k})$, $\bm{\mathbf{\Sigma_{v}}}\coloneqq$ cov$(\bm{\mathbf{v}}_{k},\bm{\mathbf{v}}_{k})$, and $\langle\bm{\mathbf{\theta}}\rangle$. (ii) $\lim_{n\to\infty}\textrm{cov}\bigr{(}\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k+n}\bigr{)}=0$. (iii) $\lim_{n\to\infty}n\cdot\textrm{cov}\bigr{(}\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k+n}-\bm{\mathbf{\theta}}_{k+n+1}\bigr{)}=0$. For example, the latter two assumptions hold true if the weight correlation function decays as $\textrm{cov}\bigr{(}\bm{\mathbf{\theta}}_{k},\bm{\mathbf{\theta}}_{k+n}\bigr{)}\propto n^{-2}$ or faster.
@@
Under the mentioned assumptions, it holds that
$\tau_{i}\coloneqq\frac{2\sigma_{\theta,i}^{2}}{\sigma_{v,i}^{2}}=\frac{\sum_{n=1}^{\infty}n\cdot{\rm cov}\bigl{(}v_{k,i},v_{k+n,i}\bigr{)}}{\sum_{n=1}^{\infty}{\rm cov}\bigl{(}v_{k,i},v_{k+n,i}\bigr{)}}$ (3)
justifying the label correlation time for this variance ratio [^1] In a continuous time setting, the definition would necessitate $\tau_{i}\coloneqq\frac{2\sigma_{\theta,i}^{2}}{\sigma_{v,i}^{2}t_{0}}$, involving a microscopic timescale $t_{0}$ to assign $\tau_{i}$ as a timescale..
Here $\sigma_{\theta,i}^{2}$ and $\sigma_{v,i}^{2}$ are eigenvalues of $\bm{\mathbf{\Sigma}}$ and $\bm{\mathbf{\Sigma_{v}}}$ respectively for a shared eigenvector [^2] The relation still holds in the case where $[\bm{\mathbf{\Sigma}},\bm{\mathbf{\Sigma_{v}}}]\neq 0$ and $\bm{\mathbf{p}}_{i}$ is not a shared eigenvector but just any vector onto which we project the dynamics. In this case, $\sigma_{\theta,i}^{2}$ and $\sigma_{v,i}^{2}$ would just be the variances of the weight and the velocity in the given direction but no longer necessarily eigenvalues of $\bm{\mathbf{\Sigma}}$ and $\bm{\mathbf{\Sigma_{v}}}$. $\bm{\mathbf{p_{i}}}$, with $\theta_{k,i}\coloneqq\vec{\theta}_{k}\cdot\vec{p_{i}}$ and $v_{k,i}\coloneqq\vec{v}_{k}\cdot\vec{p_{i}}$. The derivation is presented in Appendix D.
@@ 4.3 Variance for late training phase
In light of the autocorrelation of the noise calculated earlier, our next objective is to compute the expected weight and velocity variances at a later stage of the training. To describe the conditions of this phase, we adopt the following assumptions.
Assumption 1: Quadratic Approximation
@@
We postulate that we have reached a minimum point of the loss function, which can be adequately represented with a quadratic form as $L(\vec{\theta})=L_{0}+\frac{1}{2}(\vec{\theta}-\vec{\theta_{*}})^{\top}\bm{\mathbf{H}}(\vec{\theta}-\vec{\theta_{*}})$. We can set $L_{0}$ and $\vec{\theta_{*}}$ to zero without any loss of generality, which simplifies to:
$L(\vec{\theta})=\frac{1}{2}\vec{\theta}^{\top}\bm{\mathbf{H}}\vec{\theta}\ .$ (4)
Assumption 2: Spatially Independent Noise
We presume that the noise terms are independent of the weight vector in this late phase, represented as $\vec{\delta g}_{k}(\vec{\theta})\equiv\vec{\delta g}_{k}$.
This assumption leads to the conclusion that the noise covariance matrix is also spatially independent, $\bm{\mathbf{C}}(\vec{\theta})\equiv\bm{\mathbf{C}}$, and that the actual autocorrelation of the SGD noise now follows the relation previously calculated in Equation 2. For further motivation we refer to Appendix B.
Assumption 3: Hessian Noise Approximation
@@
We assume that the covariance of the noise commutes with the Hessian matrix, as discussed in Section 3.1,
$\left[\bm{\mathbf{C}},\bm{\mathbf{H}}\right]=0\ .$ (5)
Moreover, we assume that $10\leq\beta<$ 1 and $0<\eta\lambda_{i}<2(1+\beta)$ for all eigenvalues $\lambda_{i}$ of $\vec{H}$. If these conditions are not met, the weight fluctuations would diverge.
Calculation for one eigenvalue
With the previously stated assumptions in place, the covariance matrices $\bm{\mathbf{\Sigma}}$ and $\bm{\mathbf{\Sigma_{v}}}$ commute with $\bm{\mathbf{C}}$, $\bm{\mathbf{H}}$, and with each other (see Section C.1). As a result, they all share a common eigenbasis [^3] The relations calculated in the following paragraphs still hold in the case $\left[\bm{\mathbf{C}},\bm{\mathbf{H}}\right]\neq 0$ and where $\vec{p_{i}}$ are just the eigenvectors of the Hessian matrix. The calculated weight and velocity variances, however, are no longer eigenvalues of the corresponding covariance matrices. In Appendix M is shown what happens when $\bm{\mathbf{C}}$ and $\bm{\mathbf{H}}$ do not commute. $\vec{p_{i}}$, with $i=1,...,d$, which facilitates the computation of the expected variance. We will outline the most important steps here, while details can be found in Section C.2.
@@
For a given common eigenvector $\vec{p_{i}}$ we project the relevant variables onto this vector. This yields the projected weight $\theta_{k,i}\coloneqq\vec{p_{i}}\cdot\vec{\theta}_{k}$, velocity $v_{k,i}\coloneqq\vec{p_{i}}\cdot\vec{v}_{k}$, and noise term $\delta g_{k,i}\coloneqq\vec{p_{i}}\cdot\vec{\delta g}_{k}$ at the update step $k$. Correspondingly, we define the eigenvalues for the common eigenvector $\bm{\mathbf{p_{i}}}$ as $\lambda_{i}$ for $\vec{H}$, $\sigma_{\theta,i}^{2}$ for $\vec{\Sigma}$, $\sigma_{v,i}^{2}$ for $\vec{\Sigma_{v}}$, and $\sigma_{\delta g,i}^{2}$ for $\vec{C}$. We denote the number of batches per epoch as $M=N/S$, presuming it is an integer.
@@
By introducing the vector $\bm{\mathbf{x}}_{k,i}\coloneqq\bigl{(}\theta_{k,i}\quad\theta_{k-1,i}\bigr{)}^{\top}$ that contains not only the current weight variable but also the weight variable with a one-step time lag we can write the update equation as:
$\bm{\mathbf{x}}_{k,i}$ $=\vec{D_{i}}\vec{x}_{k-1,i}-\eta\delta g_{k,i}\vec{e}_{1}$ (6)
@@
where $\bm{\mathbf{e}}_{1}\coloneqq\bigl{(}1\quad 0\bigr{)}^{\top}$ and the matrix $\vec{D_{i}}$ governs the deterministic part of the update, its explicit expression being:
@@
$\vec{D_{i}}$ $\coloneqq\begin{pmatrix}1+\beta-\eta\lambda_{i}&-\beta\\ 1&0\end{pmatrix}\ .$ (7)
Further, we define
$\vec{E_{i}}\coloneqq\vec{D_{i}}\,{\rm cov}\left(\bm{\mathbf{x}}_{k-1,i},\delta g_{k,i}\vec{e}_{1}\right)\ .$ (8)
@@
This term encapsulates the correlation between the current weight variable and the noise term of the next update step. Typically, noise terms are assumed to be temporally uncorrelated, which would render $\vec{E_{i}}$ null. However, given the anti-correlated nature of the noise, we find a non-zero $\vec{E_{i}}$. An explicit expression of $\vec{E_{i}}$ can be found in Section C.2.
@@
With the above assumptions and definitions the following relation for the weight and velocity variances hold:
$\begin{pmatrix}\sigma_{\theta,i}^{2}\\ \sigma_{v,i}^{2}\end{pmatrix}$ $=\eta^{2}\sigma_{\delta g,i}^{2}\vec{F_{i}}\left[\vec{e}_{1}-\left(\bm{\mathbf{E_{i}}}+\vec{E_{i}}^{\top}\right)\vec{e}_{1}\right]\ ,$ (9)
where the matrix $\bm{\mathbf{F_{i}}}$ is explicitly expressed as:
$\bm{\mathbf{F_{i}}}$ $=\frac{1}{(1-\beta)\bigl{(}2(1+\beta)-\eta\lambda_{i})\bigr{)}}\begin{pmatrix}\frac{1+\beta}{\eta\lambda_{i}}&\frac{2\beta(\eta\lambda_{i}-1-\beta)}{\eta\lambda_{i}}\\ 2&2(\eta\lambda_{i}-2)\end{pmatrix}\,.$ (10)
The calculations can be found in Section C.2. The exact relation Equation 9 can be easily evaluated numerically but it can also be approximated by assuming that $M\gg 1/(1-\beta)$, which implies that the correlation time induced by momentum is substantially shorter than one epoch. Consequently, two distinct regimes of Hessian eigenvalues emerge, separated by $\lambda_{\textrm{cross}}\coloneqq 3(1-\beta)/(\eta M)$. For each of these regimes, specific simplifications apply. Notably, at $\lambda_{\textrm{cross}}$, both approximations converge.

### v3

@@ 4.1 Autocorrelation of the noise
Figure 1: Autocorrelations of the SGD noise observed over a span of 20 epochs, equivalent to 20,000 update steps. This data is collected from a later phase in the training process. The autocorrelation is projected onto 5,000 Hessian eigenvectors, and the result is averaged. The theoretical prediction Equation˜2 is also displayed along with a $2\sigma$-interval, where $\sigma$ represents the expected standard deviation of the SGD noise. The zero-point correlation is omitted as it is inherently equal to one.
We consider the correlation between two noise terms arising from different SGD update steps. Specifically, when examples are sampled without replacement while keeping the weight vector $\boldsymbol{\mathbf{\theta}}$ constant, inherent anti-correlations emerge in the noise. Below, we provide a conceptual motivation for this phenomenon and refer the reader to ˜D for a detailed derivation.
Suppose the ratio $M\coloneqq N/S$ is an integer, where $N$ is the total number of examples and $S$ is the batch size. This ensures that each epoch consists of $M$ batches, with every example presented exactly once per epoch. For a fixed weight vector $\boldsymbol{\mathbf{\theta}}$, the mean of the gradients computed over one epoch, $\boldsymbol{\mathbf{g}}_{k}(\boldsymbol{\mathbf{\theta}})$, equals the total gradient $\boldsymbol{\mathbf{\nabla}}L(\boldsymbol{\mathbf{\theta}})$. Consequently, the sum of the noise components $\boldsymbol{\mathbf{\delta g}}_{k}(\boldsymbol{\mathbf{\theta}})=\boldsymbol{\mathbf{g}}_{k}(\boldsymbol{\mathbf{\theta}})-\boldsymbol{\mathbf{\nabla}}L(\boldsymbol{\mathbf{\theta}})$ over one epoch must be zero. This implies that if a noise term points in a particular direction at the beginning of an epoch, subsequent noise terms within the same epoch are constrained to partially cancel it out, leading to anti-correlations.
Therefore, the anti-correlation between two noise terms within the same epoch can be expressed as a negative correlation proportional to the equal-time correlation, scaled by a factor of $\frac{1}{M-1}$, where $M$ is the number of batches per epoch. We assume the order of the examples within each new epoch is reshuffled, so any two noise terms from different epochs are uncorrelated. Next, we consider the probability that two batches $k$ and $k+h$, separated by $h$ update steps, belong to the same epoch. This probability is given by $\frac{M-|h|}{M}$ for $|h|\leq M$, and zero otherwise. Combining this with the anti-correlation derived above, we arrive at the correlation formula:
@@
If the total number of examples $N$ is an integer multiple of the batch size $S$ and the parameters $\boldsymbol{\mathbf{\theta}}$ of a network are kept fixed, then the autocorrelation formula for the gradient noise of an epoch-based learning schedule, where the examples for each new epoch are drawn without replacement, is given by
${\rm cov}\bigl(\vec{\delta g}_{k}(\vec{\theta}),\vec{\delta g}_{k+h}(\vec{\theta})\bigr)=\begin{cases}\boldsymbol{\mathbf{C}}(\vec{\theta})\,,&h=0\\ -\frac{M-|h|}{M(M-1)}\boldsymbol{\mathbf{C}}(\vec{\theta})\,,&1\leq|h|\leq M\\ 0\,,&|h|>M\end{cases}\ ,$ (2)
@@
The actual noise autocorrelation is illustrated in Figure˜1 and confirms the V-shaped theoretical autocorrelation structure, with the experimental details elaborated in Section˜5.2. The complete calculation is available in ˜D, highlighting the relationship $\boldsymbol{\mathbf{C}}(\vec{\theta})=(1/S)\cdot(1-S/N)\cdot\boldsymbol{\mathbf{C_{0}}}(\vec{\theta})$ with the gradient sample covariance matrix $\boldsymbol{\mathbf{C_{0}}}$.
Strictly speaking, the formula derived above applies only to a static weight vector. During training, the weights change with each update step, potentially altering this relationship. A theoretical extension accounting for fluctuating weights around a minimum is presented in ˜E, demonstrating that the predicted anti-correlations still hold approximately under realistic training dynamics in a late stage of training. Empirically, we find that in the later stages of training, the theoretical prediction given by Equation˜2 closely matches the actual training dynamics, as evidenced by the strong agreement with the data shown in Figure˜1. Further numerical investigations of the LeNet network reveal, that even at the beginning of training, when the network weights are clearly changing significantly, SGD noise still exhibits anti-correlations in good agreement with our theoretical prediction (see Section˜G.6), although our analytic theory relates only to a late phase of training.
When we sample the examples with replacement during training, there are no anti-correlations (see Section˜G.3). And while the above result assumes that $M=N/S$ is an integer, we find that the theoretical predictions remain approximately valid even when this condition is not exactly satisfied. In this case, the noise terms do not exactly add up to zero, but the variance of this sum over one epoch is reduced by a factor of $1/M$ compared to sampling with replacement. A detailed discussion of the residual noise and its effect in the non-integer case is provided in ˜F. Further numerical investigations show that in the case where $N$ is not an integer multiple of $S$, the later derived relations for the variances still hold (see Section˜G.5).
4.2 Understanding Variance Structure via Velocity Variance and Correlation Time
@@
To gain a more comprehensive understanding of the weight variance behavior, we separate it into two interpretable components: the velocity variance and a correlation time. Here, we define the velocity $\boldsymbol{\mathbf{v}}_{k}$ as the parameter update at step $k$, i.e., $\boldsymbol{\mathbf{v}}_{k}=\boldsymbol{\mathbf{\theta}}_{k+1}-\boldsymbol{\mathbf{\theta}}_{k}$. Focusing on the velocity helps to isolate how stochasticity and curvature interact dynamically. We label the ratio between weight and velocity variance, scaled by a factor of two, as the correlation time $\tau_{i}$ (see Equation˜3). This definition aligns with that of the velocity correlation time, hence justifying the nomenclature. The equivalence stands under general assumptions, such as a sufficiently fast decay of the weight correlations, $\textrm{cov}\bigr(\boldsymbol{\mathbf{\theta}}_{k},\boldsymbol{\mathbf{\theta}}_{k+n}\bigr)$, which are satisfied in our problem setup described in Section˜4.3 (see Section˜B.4), where the weight correlations will even decay exponentially fast.
To motivate this decomposition, consider the simple one-dimensional model of SGD with replacement in a quadratic loss ($L(\theta)=\frac{1}{2}\lambda\theta^{2}$) without momentum: $\theta_{k+1}=(1-\eta\lambda)\theta_{k}+\eta\delta g_{k}$}, where $\delta g_{k}$ denotes uncorrelated zero-mean gaussian gradient noise for simplicity. Assuming stationarity, it is straightforward to solve for the weight variances. For $\eta\lambda\ll 1$, this yields $\sigma_{\theta}^{2}=\frac{\eta}{2\lambda}\sigma_{\delta g}^{2}$ and exponentially decaying correlations $\langle\theta_{k+h}\theta_{k}\rangle=\sigma_{\theta}^{2}e^{-h/\tau}$ with $\tau=\frac{1}{\eta\lambda}$, since $(1-\eta\lambda)^{h}\approx e^{-\eta\lambda h}$. The velocities then exhibit anti-correlations on the same timescale, $\langle v_{k+h}v_{k}\rangle=-\frac{1}{\tau^{2}}e^{-h/\tau}\sigma_{\theta}^{2}\quad(|h|\geq 1)$. Additionally, it follows that $\sigma_{v}^{2}=\eta^{2}\sigma_{\delta g}^{2}$, and similarly to the theorem below, one recovers $\sigma_{\theta}^{2}=\frac{1}{2}\tau\sigma_{v}^{2}$. This illustrates that the velocity variance is determined by the noise, while the correlation time $\tau$ encapsulates the influence of the loss landscape curvature. For example, if the gradient noise variance scales with curvature, $\sigma_{\delta g}^{2}\propto\lambda$, the model yields constant weight variance independent of $\lambda$ – the known prediction in the with-replacement case. In the more complex case of sampling without replacement, anti-correlations in the noise modify this picture and induce a curvature-dependent suppression of $\tau$, giving rise to richer variance behavior, as explored in the next section.
[…]
@@
Under the mentioned assumptions, which are satisfied by the problem setup described in Section˜4.3, it holds that
$\tau_{i}\coloneqq\frac{2\sigma_{\theta,i}^{2}}{\sigma_{v,i}^{2}}=\frac{\sum_{n=1}^{\infty}n\cdot{\rm cov}\bigl(v_{k,i},v_{k+n,i}\bigr)}{\sum_{n=1}^{\infty}{\rm cov}\bigl(v_{k,i},v_{k+n,i}\bigr)}\ ,$ (3)
justifying the label correlation time for this variance ratio.
Here, $\sigma_{\theta,i}^{2}$ and $\sigma_{v,i}^{2}$ represent the variances of the weights and velocities, respectively, in a given direction $\boldsymbol{\mathbf{p}}_{i}$, where $\theta_{k,i}\coloneqq\boldsymbol{\mathbf{\theta}}_{k}\cdot\boldsymbol{\mathbf{p}}_{i}$ and $v_{k,i}\coloneqq\boldsymbol{\mathbf{v}}_{k}\cdot\boldsymbol{\mathbf{p}}_{i}$. If the covariance matrices $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\Sigma_{v}}}$ commute, they can be simultaneously diagonalized, allowing us to choose shared eigenvectors $\boldsymbol{\mathbf{p}}_{i}$. In this case, $\sigma_{\theta,i}^{2}$ and $\sigma_{v,i}^{2}$ become the corresponding eigenvalues. The detailed derivation is presented in ˜C.
@@ 4.3 Variance for late training phase
Given the autocorrelation of the noise calculated earlier, we aim to present the expected variances of the weights and velocities during the later stages of training. To characterize the conditions of this phase, we make the following assumptions.
@@
Assumption 1: Quadratic approximation    We assume that we have reached a minimum point of the loss function, which can be adequately represented with a quadratic form as $L(\vec{\theta})=L_{0}+\frac{1}{2}(\vec{\theta}-\vec{\theta_{*}})^{\top}\boldsymbol{\mathbf{H}}(\vec{\theta}-\vec{\theta_{*}})$. Without loss of generality, we set $L_{0}=0$ and $\vec{\theta_{*}}=0$, simplifying the expression to $L(\vec{\theta})=\frac{1}{2}\vec{\theta}^{\top}\boldsymbol{\mathbf{H}}\vec{\theta}\ .$
Assumption 2: Anti-correlated noise    We assume that the covariance of the SGD noise is static and that its autocorrelation follows the relation previously calculated in Equation˜2, even when the weight vector is not static.
While the noise generally exhibits state dependence, it is reasonable to assume a constant covariance during the later stages of training and within the time frame of our analysis. As described by Ziyin et al. [25], state dependence enters the noise through the value of the training loss. Since we observe only minimal changes in the loss during our analysis period (see Section˜G.1), assuming a static noise covariance is justified.
@@
Assumption 3: Hessian and noise covariance commute    We assume that the covariance of the noise commutes with the Hessian matrix, $\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{H}}=\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{C}}$.
This assumption is not strictly necessary, but it simplifies the analysis. The theory predicts a variance reduction in the eigenspace of the Hessian with relatively small eigenvalues without requiring commutativity (see Section˜G.7). However, when $\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{H}}\neq\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{C}}$, the calculated weight and velocity variances are no longer exact eigenvalues of their respective covariance matrices; instead, they represent variances along the directions of the Hessian eigenvectors. As long as $\boldsymbol{\mathbf{C}}$ and $\boldsymbol{\mathbf{H}}$ approximately commute – as discussed in Section˜3.1 – these variances provide a good approximation of the actual eigenvalues. For further discussion, we refer to Section˜G.7, where we demonstrate that although $\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{H}}=\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{C}}$ is not strictly satisfied in our empirical investigation, the eigenbasis of $\boldsymbol{\mathbf{H}}$ remains a good approximation for the eigenbasis of $\boldsymbol{\mathbf{C}}$. Moreover, due to finite sample sizes, the actual eigenvalues of the empirically recorded weight covariance matrix $\boldsymbol{\mathbf{\Sigma}}$ are inherently skewed. Therefore, analyzing the weight variance in the eigenbasis of $\boldsymbol{\mathbf{H}}$ is beneficial, as we explain in Section˜G.2.
Additionally, we assume that $0\leq\beta<1$ and $0<\eta\lambda_{i}<2(1+\beta)$ for all eigenvalues $\lambda_{i}$ of $\boldsymbol{\mathbf{H}}$. If these conditions are not met, the weight fluctuations would diverge. These bounds can be derived by requiring that the matrix which determines the dynamics in our model ($\boldsymbol{\mathbf{D_{i}}}$ in Equation˜4 and Equation˜5) has all eigenvalues with magnitude strictly less than one.
@@
Calculation for one eigenvalue    With the previously stated assumptions in place, the covariance matrices $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\Sigma_{v}}}$ commute with $\boldsymbol{\mathbf{C}}$, $\boldsymbol{\mathbf{H}}$, and with each other (see Section˜B.1). As a result, they all share a common eigenbasis $\vec{p_{i}}$, with $i=1,\dots,d$, which facilitates the computation of the expected variance. We will outline the most important steps here, while details can be found in Section˜B.2. For a given common eigenvector $\vec{p_{i}}$, we project the relevant variables onto this vector. This yields the projected weight $\theta_{k,i}\coloneqq\vec{p_{i}}\cdot\vec{\theta}_{k}$, velocity $v_{k,i}\coloneqq\vec{p_{i}}\cdot\vec{v}_{k}$, and noise term $\delta g_{k,i}\coloneqq\vec{p_{i}}\cdot\vec{\delta g}_{k}$ at the update step $k$. Correspondingly, we define the eigenvalues for the common eigenvector $\boldsymbol{\mathbf{p_{i}}}$ as $\lambda_{i}$ for $\vec{H}$, $\sigma_{\theta,i}^{2}$ for $\vec{\Sigma}$, $\sigma_{v,i}^{2}$ for $\vec{\Sigma_{v}}$, and $\sigma_{\delta g,i}^{2}$ for $\vec{C}$. We denote the number of batches per epoch as $M=N/S$, presuming it is an integer.
@@
By introducing the vector $\boldsymbol{\mathbf{x}}_{k,i}\coloneqq\bigl(\theta_{k,i}\quad\theta_{k-1,i}\bigr)^{\top}$ that contains not only the current weight variable but also the weight variable with a one-step time lag we can write the update equation as:
$\boldsymbol{\mathbf{x}}_{k,i}$ $=\vec{D_{i}}\vec{x}_{k-1,i}-\eta\delta g_{k,i}\vec{e}_{1}\ ,$ (4)
@@
where $\boldsymbol{\mathbf{e}}_{1}\coloneqq\bigl(1\quad 0\bigr)^{\top}$ and the matrix $\vec{D_{i}}$ governs the deterministic part of the update. Together with its explicit expression, we further define a correlation term:
@@
$\vec{D_{i}}$ $\coloneqq\begin{pmatrix}1+\beta-\eta\lambda_{i}&-\beta\\ 1&0\end{pmatrix}\ ,\ \ \ \vec{E_{i}}\coloneqq\vec{D_{i}}\,{\rm cov}\left(\boldsymbol{\mathbf{x}}_{k-1,i},\delta g_{k,i}\vec{e}_{1}\right)\ .$ (5)
@@
The term $\vec{E_{i}}$ encapsulates the correlation between the current weight variable and the noise term of the next update step. Typically, noise terms are assumed to be temporally uncorrelated, which would render $\vec{E_{i}}$ null. However, given the anti-correlated nature of the noise, we find a non-zero $\vec{E_{i}}$. An explicit expression of $\vec{E_{i}}$ can be found in Section˜B.2.
@@
With the above assumptions and definitions, the following relation for the weight and velocity variances holds:
$\begin{pmatrix}\sigma_{\theta,i}^{2}\\ \sigma_{v,i}^{2}\end{pmatrix}$ $=\eta^{2}\sigma_{\delta g,i}^{2}\vec{F_{i}}\left[\vec{e}_{1}-\left(\boldsymbol{\mathbf{E_{i}}}+\vec{E_{i}}^{\top}\right)\vec{e}_{1}\right]\ ,$ (6)
where the matrix $\boldsymbol{\mathbf{F_{i}}}$ is explicitly expressed as:
$\boldsymbol{\mathbf{F_{i}}}$ $=\frac{1}{(1-\beta)\bigl(2(1+\beta)-\eta\lambda_{i})\bigr)}\begin{pmatrix}\frac{1+\beta}{\eta\lambda_{i}}&\frac{2\beta(\eta\lambda_{i}-1-\beta)}{\eta\lambda_{i}}\\ 2&2(\eta\lambda_{i}-2)\end{pmatrix}\,$. (7)
The calculations can be found in Section˜B.2. The exact relation Equation˜6 can be easily evaluated numerically but it can also be approximated by assuming that $M\gg 1/(1-\beta)$, which implies that the correlation time induced by momentum is substantially shorter than one epoch. Consequently, two distinct regimes of Hessian eigenvalues emerge, separated by $\lambda_{\textrm{cross}}\coloneqq 3(1-\beta)/(\eta M)$. For each of these regimes, specific simplifications apply. As discussed below, the value of $\lambda_{\textrm{cross}}$ is derived as the point at which both approximations converge. As we will see, for eigenvalues below that value, the intrinsic correlation time induced by the loss landscape becomes longer than the timescale given by the anti-correlations.
Figure 2: Relationship between Hessian eigenvalues and the variances of weights and velocities, as well as correlation times. In the left panel, we present the variances of weights and velocities. The solid lines signify theoretical predictions from Equation˜6, assuming $\boldsymbol{\mathbf{C}}\approx c_{0}\boldsymbol{\mathbf{H}}$ with $c_{0}$ fitted to the data. The right panel showcases the correlation time together with the theoretical prediction resulting from Equation˜6, which does not require the $\boldsymbol{\mathbf{C}}\approx c_{0}\boldsymbol{\mathbf{H}}$ assumption.